### v1

@@ 4. Application for two–dimensional materials
In this section, we apply the previous theory to the case of two–dimensional materials in the tight–binding approximation. Our goal is to describe the spectrum of such materials in the presence of a soft-wall. In the one-dimensional case, we could without loss of generality fix the unit cell to be $[0,1)$ and the corresponding Fourier variable belongs to $[-\pi,\pi)$. To describe general two-dimensional models, we need the full terminology of solid state physics.
@@ 4.1.1. Bravais lattices and bulk operators in tight-binding models
$\forall\Psi\in\ell^{2}(\mathbb{L}_{1},\mathbb{C}^{M}),\quad\forall{\mathbf{R}}_{1}\in\mathbb{L}_{1}\qquad\left(H_{{\mathbf{k}}_{2}}\Psi\right)_{{\mathbf{R}}_{1}}=\sum_{{\mathbf{R}}_{1}^{\prime}\in\mathbb{L}_{1}}h_{{\mathbf{k}}_{2}}({\mathbf{R}}_{1}^{\prime})\Psi_{{\mathbf{R}}_{1}-{\mathbf{R}}_{1}^{\prime}},$
@@
In what follows, we fix ${\mathbf{k}}_{2}\in\Gamma_{2}^{*}$, and study the one-dimensional periodic operator $H_{{\mathbf{k}}_{2}}$, and its soft–wall counterpart $H_{{\mathbf{k}}_{2}}^{\sharp}(t)$. The spectrum of $H_{{\mathbf{k}}_{2}}$ is purely essential, composed of $M$ (one-dimensional) Bloch bands. It is the union of the spectra of $\left(H_{{\mathbf{k}}_{2}}\right)_{{\mathbf{k}}_{1}}$, defined as (compare with (4))
@@ 4.1.2. Soft walls parallel to a lattice vector
where $w_{*}:\mathbb{R}\mapsto{\mathcal{S}}_{M}(\mathbb{C})$ is a one–dimensional Lipschitz soft wall potential as in Section 2.1.2. Note that $w({\mathbf{x}}+\lambda{\mathbf{a}}_{2})=w({\mathbf{x}})$, so this wall is constant along the ${\mathbf{a}}_{2}$ direction. In addition, since ${\mathbf{a}}_{2}^{\perp}$ is normalized, this potential has the same Lipschitz constant as $w_{1}$. To introduce a dislocation parameter $t\in\mathbb{R}$, we set
@@
Finally, the soft-wall model is the operator
@@
We recover the soft-wall described in Section 2.1.2: the normalization for $t$ in (21) has been chosen so that when $t$ moves from 00 to $1$, the wall shifts by the vector ${\mathbf{a}}_{1}$. We can now apply the previous results to this case.
@@ 4.1.3. Edge spectrum in the two–dimensional setting
Assume that $w_{1}:\mathbb{R}\mapsto{\mathcal{S}}_{M}$ is a Lipschitz soft–wall. Then, with the previous notation,
@@
    ${\mathcal{D}}:=\left\{\Psi\in\ell^{2}(\mathbb{L}_{1},\mathbb{C}^{M}),\qquad W_{t=0}\Psi\in\ell^{2}(\mathbb{L}_{1},\mathbb{C}^{M})\right\}.$
@@
    The map $\Gamma^{*}_{2}\times\mathbb{R}\ni({\mathbf{k}}_{2},t)\mapsto H^{\sharp}_{{\mathbf{k}}_{2}}(t)$ is norm-resolvent continuous and $1$-quasi–periodic in $t$.
@@
Let $w:\mathbb{R}\mapsto{\mathcal{S}}_{N}$ be a soft–wall which is $\nu$–Lipschitz, for some $\nu>0$. Then, for all ${\mathbf{k}}_{2}\in\Gamma^{*}_{2}$, all energy $E$ belonging to an essential gap of width $g(E)\geq\nu|\Gamma|\|{\mathbf{a}}_{2}\|^{-1}$ of $H_{{\mathbf{k}}_{2}}$, and all $t_{0}\in\mathbb{R}$, the operator $H^{\sharp}_{{\mathbf{k}}_{2}}(t_{0})$ has at least ${\mathcal{N}}(E)$ eigenvalues in each interval of the form $(\lambda,\lambda+\nu|\Gamma|\|{\mathbf{a}}_{2}\|^{-1}]$ in this gap.
@@
We skip its proof, as it is similar to the one of Theorem 1.3. This time, the density of eigenvalues is of order $\frac{\nu|\Gamma|}{{\mathcal{N}}_{{\mathbf{k}}_{2}}(E)\|{\mathbf{a}}_{2}\|}$.
@@ 4.2. Walls in other directions
In this section, the wall will be aligned with the vector $\widetilde{{\mathbf{a}}}_{2}$, and move along the $\widetilde{{\mathbf{a}}}_{1}$–direction (which is also the ${\mathbf{a}}_{1}$–direction). Varying $n$ and $m$ allows to describe all commensurable directions for the wall, expect for the ${\mathbf{a}}_{1}$ and ${\mathbf{a}}_{2}$–directions (which were studied in the previous section).
@@ 4.2.1. Bulk spectrum
$\bigl{(}\widetilde{\mathsf{H}}\Phi\bigr{)}({\mathbf{R}})=\sum_{{\mathbf{R}}^{\prime}\in\widetilde{\mathbb{L}}}\widetilde{\mathsf{h}}({\mathbf{R}}^{\prime})\Phi_{{\mathbf{R}}-{\mathbf{R}}^{\prime}}\quad\text{ where }\widetilde{\mathsf{h}}({\mathbf{R}}^{\prime})_{ij}={\mathsf{h}}({\mathbf{R}}^{\prime}-{\mathbf{y}}_{i}+{\mathbf{y}}_{j}).$
@@
$u_{{\mathbf{K}}}:=\left({\mathrm{e}}^{-{\mathrm{i}}{\mathbf{y}}_{1}\cdot({\mathbf{k}}+{\mathbf{K}})}u,\cdots,{\mathrm{e}}^{-{\mathrm{i}}{\mathbf{y}}_{L}\cdot({\mathbf{k}}+{\mathbf{K}})}u\right)^{\top}$
@@
$\sigma\left(\widetilde{H}_{\widetilde{{\mathbf{k}}}_{2}}\right)=\bigcup_{\widetilde{{\mathbf{k}}}_{1}\in\widetilde{\Gamma}_{1}^{*}}\sigma\left(\widetilde{{\mathsf{H}}}(\widetilde{{\mathbf{k}}}_{1}+\widetilde{{\mathbf{k}}}_{2})\right)$
@@
It remains to prove that the union in $\widetilde{{\mathbf{K}}}_{2}^{*}$ can be dropped. Recall that we assume that $n$ and $m$ are coprime. Let $p,q\in\mathbb{Z}^{2}$ so that $pn+qm=1$. We have, using (24),
@@ 4.2.2. Spectral flow for the edge Hamiltonian with general angles
Let $w:\mathbb{R}\mapsto{\mathcal{S}}_{N}$ be a soft–wall which is $\nu$–Lipschitz, for some $\nu>0$. If $E$ belongs to an essential gap of width $g(E)\geq\nu$, then for all $t_{0}\in\mathbb{R}$ the operator $\widetilde{H}^{\sharp}(t_{0})_{\widetilde{{\mathbf{k}}}_{2}}$ has at least ${\mathcal{N}}(E)$ eigenvalues in each interval of the form $(\lambda,\lambda+\nu|\Gamma|/\|\widetilde{\mathbf{a}}_{2}\|]$ in this gap.
Note that the density of edge states is of order ${\mathcal{N}}(E)\|\widetilde{\mathbf{a}}_{2}\|/(\nu|\Gamma|)$. When the cut becomes close to incommensurate, in the sense that $\|\widetilde{\mathbf{a}}_{2}\|$ becomes large, then there are numerous edge states for each value of $\widetilde{{\mathbf{k}}}_{2}$.
@@ 4.3.1. The Wallace model
Let us start by describing the bulk spectrum of graphene. Graphene has a honeycomb lattice structure, which is the triangular Bravais lattice, with two atoms per unit cell. As basis vectors, we may chose
@@
${\mathbf{a}}_{1}^{*}=\frac{2\pi}{\sqrt{3}a_{0}}\begin{pmatrix}\sqrt{3}\\ -1\end{pmatrix},\qquad{\mathbf{a}}_{2}^{*}=\frac{2\pi}{\sqrt{3}a_{0}}\begin{pmatrix}\sqrt{3}\\ 1\end{pmatrix}$
@@ 4.3.2. Zigzag orientation
The gap of the one-dimensional operator $H_{{\mathbf{k}}_{2}}$ closes only for ${\mathbf{k}}_{2}=\frac{1}{3}{\mathbf{a}}_{2}^{*}+\mathbb{L}_{2}^{*}$ and ${\mathbf{k}}_{2}=\frac{2}{3}{\mathbf{a}}_{2}^{*}+\mathbb{L}_{2}^{*}$, corresponding to the projections of the ${\mathbf{K}}$ and ${\mathbf{K}}^{\prime}$ points on the line ${\mathbf{a}}_{2}^{*}\mathbb{R}$, along the ${\mathbf{a}}_{1}^{*}$–direction. This can be seen in the first panel of Figure 5. Apart from these two points, we have ${\mathcal{N}}_{{\mathbf{k}}_{2}}(E=0)=1$. Numerical simulations of the edge modes are presented in Figures 6 and 7.
@@ 4.3.4. Another rational cut
$f({\mathbf{0}})=\begin{pmatrix}1&0\\ 1&1\end{pmatrix},\qquad f(-\overline{{\mathbf{a}}}_{1})=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\qquad f(\overline{{\mathbf{a}}}_{1}-\overline{{\mathbf{a}}}_{2})=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$
@@ 4.3.5. Numerical illustration of the edge spectrum for different cuts
The corresponding edge models ${\mathsf{H}}^{\sharp}(t)$ (zigzag orientation), $\widetilde{{\mathsf{H}}}^{\sharp}(t)$ (armchair orientation) and $\overline{{\mathsf{H}}}^{\sharp}(t)$ (our commensurate angle with $n=2$ and $m=-1$, called angle in what follows) are $1$–quasi–periodic. Recall that our results shows that, outside the projection of the Dirac cones, we must have
@@
Acknowledgements
The research leading to these results has received funding from CNRS, AAP IRL 2022, and from ANID, Chile, through the CMM Basal grant FB210005 and through Fondecyt Project 11220194.
@@ A.1. Norm resolvent topology
where we set ${\mathcal{R}}(z,A):=(z-A)^{-1}$. This endows ${\mathcal{A}}({\mathcal{H}})$ with the norm-resolvent topology. Recall that the point ${\mathrm{i}}\in\mathbb{C}$ does not play any special role, and that, for all $z\in\mathbb{C}\setminus\mathbb{R}$, the corresponding metric with $z$ instead of ${\mathrm{i}}$ defines the same topology. More specifically, if $d_{z}(A,B):=\|(z-A)^{-1}-(z-B)^{-1}\|_{\rm op}$, we have the following.
@@
Proof.
We follow [23, Remark IV.3.13]. For $z,z_{0}\in\mathbb{C}$, we formally have
${\mathcal{R}}(z,A)-{\mathcal{R}}(z,B)=\frac{z_{0}-A}{z-A}\left[{\mathcal{R}}(z_{0},A)-{\mathcal{R}}(z_{0},B\right]\frac{z_{0}-B}{z-B}.$
In addition, since $\frac{z-B}{z_{0}-B}=1+(z-z_{0}){\mathcal{R}}(z_{0},B)$ and, similarly, $\frac{z-A}{z_{0}-A}=1+(z-z_{0}){\mathcal{R}}(z_{0},A)$, we deduce that
$\frac{z-B}{z_{0}-B}$ $=\frac{z-A}{z_{0}-A}+(z-z_{0})\left[{\mathcal{R}}(z_{0},B)-{\mathcal{R}}(z_{0},A)\right]$
$=\frac{z-A}{z_{0}-A}\left[1+(z-z_{0})\frac{z_{0}-A}{z-A}\left[{\mathcal{R}}(z_{0},B)-{\mathcal{R}}(z_{0},A)\right]\right].$
In particular, if $(z_{0}-A)$ and $(z-A)$ are invertible with bounded inverse, and if $\|{\mathcal{R}}(z_{0},B)-{\mathcal{R}}(z_{0},A)\|_{\rm{op}}$ is sufficiently small, then we deduce that $(z-B)$ is also invertible. Actually,
${\mathcal{R}}(z,A)-{\mathcal{R}}(z,B)=\frac{z_{0}-A}{z-A}\left[{\mathcal{R}}(z_{0},A)-{\mathcal{R}}(z_{0},B\right]\left[1+(z-z_{0})\frac{z_{0}-A}{z-A}\left[{\mathcal{R}}(z_{0},B)-{\mathcal{R}}(z_{0},A)\right]\right]^{-1}\frac{z_{0}-A}{z-A}.$
In particular, with $z_{0}={\mathrm{i}}$,
$d_{z}(A,B)\leq\frac{\left\|\frac{{\mathrm{i}}-A}{z-A}\right\|_{\rm{op}}^{2}d(A,B)}{1-|z-{\mathrm{i}}|\left\|\frac{{\mathrm{i}}-A}{z-A}\right\|_{\rm{op}}d(A,B)},$
which makes sense whenever $d(A,B)<\left(|z-{\mathrm{i}}|\|\frac{{\mathrm{i}}-A}{z-A}\|_{{\rm{op}}}\right)^{-1}$. ∎

### v3

@@ 4. Application for two–dimensional materials
In this section, we apply the previous theory to the case of two–dimensional materials in the tight–binding approximation. Our goal is to describe the spectrum of such materials in the presence of a soft wall. In the one-dimensional case, we could without loss of generality fix the unit cell to be $[0,1)$ and the corresponding Fourier variable belongs to $[-\pi,\pi)$. To describe general two-dimensional models, we need the full terminology of solid state physics.
@@ 4.1.1. Bravais lattices and bulk operators in tight-binding models
$\forall\Psi\in\ell^{2}(\mathbb{L}_{1},\mathbb{C}^{M}),\quad\forall{\mathbf{R}}_{1}\in\mathbb{L}_{1},\qquad\left(H_{{\mathbf{k}}_{2}}\Psi\right)_{{\mathbf{R}}_{1}}=\sum_{{\mathbf{R}}_{1}^{\prime}\in\mathbb{L}_{1}}h_{{\mathbf{k}}_{2}}({\mathbf{R}}_{1}^{\prime})\Psi_{{\mathbf{R}}_{1}-{\mathbf{R}}_{1}^{\prime}},$
@@
In what follows, we fix ${\mathbf{k}}_{2}\in\Gamma_{2}^{*}$, and study the one-dimensional periodic operator $H_{{\mathbf{k}}_{2}}$, and its soft wall counterpart $H_{{\mathbf{k}}_{2}}^{\sharp}(t)$. The spectrum of $H_{{\mathbf{k}}_{2}}$ is purely essential, composed of $M$ (one-dimensional) Bloch bands. It is the union of the spectra of $\left(H_{{\mathbf{k}}_{2}}\right)_{{\mathbf{k}}_{1}}$, defined as (compare with (5))
@@ 4.1.2. Soft walls parallel to a lattice vector
where $w_{*}:\mathbb{R}\mapsto{\mathcal{S}}_{M}(\mathbb{C})$ is a one–dimensional Lipschitz soft wall potential as in Section 2.1.2. Note that $w({\mathbf{x}}+\lambda{\mathbf{a}}_{2})=w({\mathbf{x}})$, so this wall is constant along the ${\mathbf{a}}_{2}$ direction. In addition, since ${\mathbf{a}}_{2}^{\perp}$ is normalized, this potential has the same Lipschitz constant as $w_{1}$. To introduce a dislocation parameter $t\in\mathbb{R}$, we set (see also Remark 1.4)
@@
Finally, the soft wall model is the operator
@@
We recover the soft wall described in Section 2.1.2: the normalization for $t$ in (24) has been chosen so that when $t$ moves from 00 to $1$, the wall shifts by the vector ${\mathbf{a}}_{1}$. We can now apply the previous results to this case.
@@ 4.1.3. Edge spectrum in the two–dimensional setting
Assume that $w_{1}:\mathbb{R}\mapsto{\mathcal{S}}_{M}$ is a Lipschitz soft wall. Then, with the previous notation,
@@
    ${\mathcal{D}}:=\left\{\Psi\in\ell^{2}(\mathbb{L}_{1},\mathbb{C}^{M}),\quad W_{t=0}\Psi\in\ell^{2}(\mathbb{L}_{1},\mathbb{C}^{M})\right\}.$
@@
    The map $\Gamma^{*}_{2}\times\mathbb{R}\ni({\mathbf{k}}_{2},t)\mapsto H^{\sharp}_{{\mathbf{k}}_{2}}(t)$ is norm-resolvent continuous and translation equivariant in $t$.
@@
Let $w:\mathbb{R}\mapsto{\mathcal{S}}_{N}$ be a soft wall which is $\nu$–Lipschitz, for some $\nu>0$. Then, for all ${\mathbf{k}}_{2}\in\Gamma^{*}_{2}$, all energy $E$ belonging to an essential gap of width $g(E)\geq\nu|\Gamma|\|{\mathbf{a}}_{2}\|^{-1}$ of $H_{{\mathbf{k}}_{2}}$, and all $t_{0}\in\mathbb{R}$, the operator $H^{\sharp}_{{\mathbf{k}}_{2}}(t_{0})$ has at least ${\mathcal{N}}(E)$ eigenvalues in each interval of the form $(\lambda,\lambda+\nu|\Gamma|\|{\mathbf{a}}_{2}\|^{-1}]$ in this gap.
@@
We skip its proof, as it is similar to the one of Theorem 1.3 (see also Remark 1.4). This time, the density of eigenvalues is lower bounded by $\frac{\nu|\Gamma|}{{\mathcal{N}}_{{\mathbf{k}}_{2}}(E)\|{\mathbf{a}}_{2}\|}$.
@@ 4.2. Walls in other directions
In this section, the wall will be aligned with the vector $\widetilde{{\mathbf{a}}}_{2}$, and move along the $\widetilde{{\mathbf{a}}}_{1}$–direction (which is also the ${\mathbf{a}}_{1}$–direction). Varying $n$ and $m$ allows to describe all commensurable directions for the wall, except for the ${\mathbf{a}}_{1}$ and ${\mathbf{a}}_{2}$–directions (which were studied in the previous section).
@@ 4.2.1. Bulk spectrum
$\bigl{(}\widetilde{\mathsf{H}}\Phi\bigr{)}({\mathbf{R}})=\sum_{{\mathbf{R}}^{\prime}\in\widetilde{\mathbb{L}}}\widetilde{\mathsf{h}}({\mathbf{R}}^{\prime})\Phi_{{\mathbf{R}}-{\mathbf{R}}^{\prime}}\quad\text{ where }\quad\widetilde{\mathsf{h}}({\mathbf{R}}^{\prime})_{ij}={\mathsf{h}}({\mathbf{R}}^{\prime}-{\mathbf{y}}_{i}+{\mathbf{y}}_{j}).$
@@
$u_{{\mathbf{K}}}:=\left({\mathrm{e}}^{-{\mathrm{i}}{\mathbf{y}}_{1}\cdot({\mathbf{k}}+{\mathbf{K}})}u,\cdots,{\mathrm{e}}^{-{\mathrm{i}}{\mathbf{y}}_{L}\cdot({\mathbf{k}}+{\mathbf{K}})}u\right)^{\top},$
@@
$\sigma\left(\widetilde{H}_{\widetilde{{\mathbf{k}}}_{2}}\right)=\bigcup_{\widetilde{{\mathbf{k}}}_{1}\in\widetilde{\Gamma}_{1}^{*}}\sigma\left(\widetilde{{\mathsf{H}}}(\widetilde{{\mathbf{k}}}_{1}+\widetilde{{\mathbf{k}}}_{2})\right).$
@@
It remains to prove that the union in $\widetilde{{\mathbf{K}}}_{2}^{*}$ can be dropped. Recall that we assume that $n$ and $m$ are coprime. Let $(p,q)\in\mathbb{Z}^{2}$ so that $pn+qm=1$. We have, using (27),
@@ 4.2.2. Spectral flow for the edge Hamiltonian with general angles
Let $w:\mathbb{R}\mapsto{\mathcal{S}}_{N}$ be a soft wall which is $\nu$–Lipschitz, for some $\nu>0$. If $E$ belongs to an essential gap of width $g(E)\geq\nu$, then for all $t_{0}\in\mathbb{R}$ the operator $\widetilde{H}^{\sharp}(t_{0})_{\widetilde{{\mathbf{k}}}_{2}}$ has at least ${\mathcal{N}}(E)$ eigenvalues in each interval of the form $(\lambda,\lambda+\nu|\Gamma|/\|\widetilde{\mathbf{a}}_{2}\|]$ in this gap.
Note that the density of edge states is lower bounded by ${\mathcal{N}}(E)\|\widetilde{\mathbf{a}}_{2}\|/(\nu|\Gamma|)$. When the cut becomes close to incommensurate, in the sense that $\|\widetilde{\mathbf{a}}_{2}\|$ becomes large, then there are numerous edge states for each value of $\widetilde{{\mathbf{k}}}_{2}$.
@@ 4.3.1. The Wallace model
Let us start by describing the bulk spectrum of graphene. Graphene has a honeycomb lattice structure, which is the triangular Bravais lattice, with two atoms per unit cell. As basis vectors, we may choose
@@
${\mathbf{a}}_{1}^{*}=\frac{2\pi}{\sqrt{3}a_{0}}\begin{pmatrix}\sqrt{3}\\ -1\end{pmatrix},\qquad{\mathbf{a}}_{2}^{*}=\frac{2\pi}{\sqrt{3}a_{0}}\begin{pmatrix}\sqrt{3}\\ 1\end{pmatrix}.$
@@ 4.3.2. Zigzag orientation
The gap of the one-dimensional operator $H_{{\mathbf{k}}_{2}}$ closes only for ${\mathbf{k}}_{2}=\frac{1}{3}{\mathbf{a}}_{2}^{*}+\mathbb{L}_{2}^{*}$ and ${\mathbf{k}}_{2}=\frac{2}{3}{\mathbf{a}}_{2}^{*}+\mathbb{L}_{2}^{*}$, corresponding to the projections of the ${\mathbf{K}}$ and ${\mathbf{K}}^{\prime}$ points on the line ${\mathbf{a}}_{2}^{*}\mathbb{R}$, along the ${\mathbf{a}}_{1}^{*}$–direction. This can be seen in the first panel of Figure 6. Apart from these two points, we have ${\mathcal{N}}_{{\mathbf{k}}_{2}}(E=0)=1$. Numerical simulations of the edge modes are presented in Figures 7 and 8.
@@ 4.3.4. Another rational cut
$f({\mathbf{0}})=\begin{pmatrix}1&0\\ 1&1\end{pmatrix},\qquad f(-\overline{{\mathbf{a}}}_{1})=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\qquad f(\overline{{\mathbf{a}}}_{1}-\overline{{\mathbf{a}}}_{2})=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},$
@@ 4.3.5. Numerical illustration of the edge spectrum for different cuts
The corresponding edge models ${\mathsf{H}}^{\sharp}(t)$ (zigzag orientation), $\widetilde{{\mathsf{H}}}^{\sharp}(t)$ (armchair orientation) and $\overline{{\mathsf{H}}}^{\sharp}(t)$ (our commensurate angle with $n=2$ and $m=-1$, called angle in what follows) are translation equivariant. Recall that our results shows that, outside the projection of the Dirac cones, we must have
@@
Similar spectral flows have been observed in a dislocated model for wave propagation in hexagonal structures, see [7].
@@ A.1. Norm resolvent topology
where we set ${\mathcal{R}}(z,A):=(z-A)^{-1}$. This endows ${\mathcal{A}}({\mathcal{H}})$ with the norm-resolvent topology. Recall that the point ${\mathrm{i}}\in\mathbb{C}$ does not play any special role, and that, for all $z\in\mathbb{C}\setminus\mathbb{R}$, the corresponding metric with $z$ instead of ${\mathrm{i}}$ defines the same topology. More specifically, we have the following.
@@
We refer to [27, Remark IV.3.13] for the proof, see in particular [27, Eqn. IV (3.10)].